# Jason Gaitonde

### v1

@@ Abstract
A popular method for sampling from high-dimensional distributions is the Gibbs sampler, which iteratively resamples sites from the conditional distribution of the desired measure given the values of the other coordinates. But to what extent does the order of site updates matter in the mixing time? Two natural choices are (i) standard, or random scan, Glauber dynamics where the updated variable is chosen uniformly at random, and (ii) the systematic scan dynamics where variables are updated in a fixed, cyclic order. We first show that for systems of dimension $n$, one round of the systematic scan dynamics has spectral gap at most a factor of order $n$ worse than the corresponding spectral gap of a single step of Glauber dynamics, tightening existing bounds in the literature by He, et al. [NeurIPS ’16] and Chlebicka, Łatuszyński, and Miasodejow [Ann. Appl. Probab. ’24]. This result is sharp even for simple spin systems by an example of Roberts and Rosenthal [Int. J. Statist. Prob. ’15]. We complement this with a converse statement: if all, or even just one scan order rapidly mixes, the Glauber dynamics has a polynomially related mixing time, resolving a question of Chlebicka, Łatuszyński, and Miasodejow. Our arguments are simple and only use elementary linear algebra and probability.
1 Introduction and main results
@@
An important task in statistics, computer science, and machine learning is sampling from high-dimensional distributions. However, the so-called “curse of dimensionality” can make this problem provably computationally hard in general [Sly10, SS12, GKK24]; while it is often the case that determining the probability of a configuration up to a universal normalizing constant (also called the partition function) is easy, the exact computation of this normalization requires summing over (at least) exponentially many configurations in the dimension. Therefore, a popular option in both theory and practice is to run a suitable ergodic Markov chain on the state space that is easy to implement and whose stationary measure is the desired distribution.
@@
Of these methods, an extremely popular approach for sampling from high-dimensional distributions is the well-known Gibbs sampler introduced by Geman and Geman [GG84] inspired by ideas from statistical physics. The Gibbs sampler is particularly convenient since it is usually quite efficient to implement and may rapidly mix towards the stationary distribution (i.e. in only polynomial number of steps in the dimension). Concretely, the general form of these procedures is as follows: for some target distribution $\pi$ on a product space $\mathcal{X}:=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$ that one wishes to sample from, consider any starting configuration $X^{0}\in\mathcal{X}$. At each time $t\geq 1$, the sampler chooses a coordinate $i_{t}\in[n]$ according to some predetermined (possibly randomized) rule and rerandomizes the $i_{t}$’th coordinate of $X^{t-1}$ according to the distribution of $\pi$ on $\mathcal{X}_{i_{t}}$ conditioned on the remaining coordinates agreeing with $X^{t-1}$. All other coordinates remain the same. It is not difficult to show that $\pi$ is stationary with respect to these dynamics for any choice of indices $i_{t}$. Evaluating these conditional distributions is often computationally straightforward, so the overall complexity of the method is governed by the number of steps for which it needs to run to be close to stationarity.
@@
Two natural ways to choose the coordinate $i_{t}$ to update at time $t$ are (i) standard (random scan) Glauber dynamics, where $i_{t}\in[n]$ is chosen uniformly at random, and (ii) the systematic scan, where $i_{t}=\sigma(t\ (\mathrm{mod}\ n))$ for some fixed choice of permutation $\sigma:[n]\to[n]$ so that the choice of sites to update cycles around the coordinates in some fixed order. For any choice of updates, the crucial quantity of interest is the mixing time to stationarity: how many steps of the Gibbs sampler are needed until the distribution at time $T$ is close to $\pi$ in, say, total variation distance? Ideally, the process mixes in time $\mathsf{poly}(n)$ or even more ambitiously $O(n\log n)$, so that the overall sampling complexity scales only polynomially in the dimension.
@@
By now, a variety of advanced methods have been introduced to bound the mixing time of important Markov chains (see, e.g. [LP17, MT06] for textbook treatments). For high-dimensional distributions on product spaces, mixing time analyses are typically studied for the Glauber dynamics since this Markov chain enjoys several analytical advantages. For one, the Markov chain is reversible with respect to $\pi$ so that the mixing properties can be analyzed by well-understood spectral methods. In particular, the mixing time connects naturally to linear-algebraic machinery via the spectral theorem, making these dynamics significantly more amenable to a rigorous study in most applications. Moreover, the Glauber dynamics treats sites symmetrically and is the natural tensorized analogue of the transition operators that operates independently on each coordinate, as would be the case for product distributions.
@@
By contrast, the transition operator associated with the systematic scan is importantly not typically reversible once $n>2$ except in simple settings, like when $\pi$ is a product measure. Even though the transitions are a product of self-adjoint operators, the lack of commutativity makes the analysis highly nontrivial. In fact, this issue is not just a feature of existing technical machinery: the actual site ordering can indeed significantly affect the mixing time since the application of the transitions varies deterministically depending on the iteration [RR15, HSMR16]. Nonetheless, from the computational side, the fixed and sequential nature of the systematic scan is among several practical considerations that can affect runtime in real-world implementations [ZR13]. No additional randomness is needed in the sequence of updated sites and moreover, one can exploit locality among adjacent sites in the updating process, which can yield practical gains given hardware considerations. For these reasons, experimental evaluations are often done using the systematic scan [DGJ06b].
@@
As a result, it is of significant importance to understand the relative performance between these two implementations of the Gibbs sampler. More generally, to what extent does the choice of site updates affect the mixing time of important high-dimensional distributions one wishes to sample from? From an analytical perspective, it may be significantly easier to prove mixing time bounds for Glauber dynamics for the reasons above; but then to what extent can these guarantees extend to the systematic scan? Obtaining (hopefully sharp) black-box guarantees of this type provide concrete performance guarantees for the systematic scan.
@@
Understanding which deterministic sequences of site updates mix can also be of importance in applications beyond sampling, like in statistical applications where one does not have perfect knowledge of how sites are chosen to update. For instance, the Glauber dynamics are also considered in economics as “noisy best-response dynamics” for decentralized game dynamics [KMR93, Blu93, You11] as well as in statistical physics as a model for a particle system to converge towards thermal equilibrium. In these systems, one views the dynamics as arising exogenously and instead wishes to learn some macroscopic properties or statistics of typical samples from $\pi$. However, it may not be possible to know the precise (possibly stochastic) process that generates site updates; one can instead only observe the sequence of site updates of the system of interest, so one cannot quite be sure that the system has truly mixed near $\pi$ even if they know that Glauber dynamics mixes. Can a statistician nonetheless conclude that after some predictable period of time, the observed configuration is close to the true stationary measure without knowledge of the randomness in the sequence of site updates? This amounts to asserting that the configuration has conditional law close to $\pi$ even given the sequence of updated sites (so long as this sequence was chosen independently of the realization of the dynamics), which is a stronger notion of mixing.
In this paper, we answer these questions positively. We first prove the following result on mixing times. To state the result, let $t^{\mathsf{GD}}_{\mathrm{mix}}(\varepsilon)$ denote the time for Glauber dynamics to be within $\varepsilon$ of $\pi$ in total variation distance from any starting configuration, and let $t^{\mathsf{SS}(\sigma)}_{\mathrm{mix}}(\varepsilon)$ denote the time for Glauber dynamics to be within $\varepsilon$ of $\pi$ in total variation distance from any starting configuration, where the scan is done in order $\sigma:[n]\to[n]$ (see Section 2 for more formal definitions).
@@
Since the systematic scan consists of $n$ site updates, Theorem 1.1 asserts that the mixing time of the systematic scan is no more than a factor $n^{2}$ worse than that of Glauber dynamics up to the logarithmic factor inherent to spectral analyses of mixing. Importantly, Theorem 1.1 is not improvable in the worst-case even in simple (anti-ferromagnetic) spin systems as witnessed by an example of Roberts and Rosenthal [RR15]. For completeness, we provide a simple, self-contained proof of this example following their reasoning in Section 4. Thus, Theorem 1.1 constitutes the sharpest possible upper bound on the mixing time of the systematic scan that can hold unconditionally for all systematic scans for Gibbs samplers in terms of the mixing time of Glauber dynamics, at least up to the unavoidable logarithmic factor.
We further show that a version of Theorem 1.1 holds for almost every pseudorandom sequence of sites to update. When translated to the Glauber dynamics, or most similar dynamics, our argument yields an efficient certification procedure that a process has mixed (or at least contracted in total variation) even conditioned on the site updates that holds with $1-o(1)$ probability on a random sequence of appropriate length from Glauber. Thus, in economics and statistical physics applications, one can formally assert that the samples one obtains approximate $\pi$ after an amount of time that depends only the spectral gap of Glauber, even if one does not know whether the actual process exactly followed the Glauber dynamics. Any sufficiently pseudorandom sequence of site updates will do, regardless of how they are generated (so long as it is independent of the actual updates), and this condition is easy to certify.
@@
The proof of Theorem 1.1 and the related applications is derived from purely linear-algebraic arguments that bound the operator norms of appropriate matrices. The connection between spectral gaps and mixing times is fairly standard, so we only briefly remark on this connection here and defer the standard proofs to Section 2. At a high level, it is well-known that the Glauber dynamics having relaxation time to stationarity at most $1/\delta$ is the same as asserting that for any function $f\in L^{2}(\mathcal{X},\pi)$ with $\text{Var}_{\pi}(f)=1$ and $\mathbb{E}_{\pi}[f]=0$,
@@
for appropriate linear operators $\mathsf{P}_{i}$ that are self-adjoint with respect to the natural inner product associated to $L^{2}(\mathcal{X},\pi)$. These operators act on functions $f\in L^{2}(\mathcal{X},\pi)$ by taking the conditional expectation of $f$ after rerandomizing the $i$th site given the remaining coordinates. Our first main technical result establishes a suitable operator norm bound on the product of transition matrices associated to any systematic scan. By standard arguments relating mixing times and operator norms even for non-reversible Markov chains, we obtain Theorem 1.1.
In fact, this argument can be more or less reversed up to polynomial factors. We further show that if any systematic scan has a spectral gap, then Glauber dynamics has a polynomially related spectral gap, which therefore implies a mixing bound by standard results. The spectral gap comparison resolves Open Problem 3.7 of Chlebicka, Łatuszyński, and Miasodejow [CŁM24]. In fact, the following polynomial comparison of mixing times holds:
Theorem 1.2.
The following holds for any $0<\varepsilon\leq 1$. Let $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$ be a product of finite state spaces and let $\pi$ be any distribution on $\mathcal{X}$. Then
We provide the precise quantitative dependencies in Section 3; the bound improves depending on the assumption made on whether all systematic scan orders $\sigma$ mix fast, or whether there is a single one that mixes fast. While these bounds are significantly weaker than our tight results for the converse direction, Theorem 1.1 and Theorem 1.2 together constitute an unconditional polynomial equivalence between the mixing times of the Glauber dynamics and any systematic scan. The mixing time comparison follows from the spectral gap comparison by a recent comparison of the mixing time of non-reversible chains in terms of the minimal nontrivial singular value of the Laplacian by Chatterjee [Cha23]. While this value is usually not related to the operator norm for general non-reversible Markov chains, the projection property of the Gibbs sampler will enable us to relate these quantities up to polynomial factors.
@@
Organization. In Section 1.1, we discuss related bounds and comparisons in the literature between Glauber dynamics and the systematic scan; to our knowledge, our spectral gap comparison subsumes all such existing bounds and cannot be improved in light of known lower bounds. In Section 2, we include for completeness standard results on the relation between the spectral gap and mixing times, including for non-reversible chains, for our main application. In Section 3, we prove the requisite linear-algebraic facts that yield our mixing time comparisons. Finally, we show in Section 4 that Theorem 1.1 cannot be significantly improved via simple, existing examples.

### v2

@@ Abstract
A popular method for sampling from high-dimensional distributions is the Gibbs sampler, which iteratively resamples sites from the conditional distribution of the desired measure given the values of the other coordinates. It is natural to ask to what extent does the order of site updates matter in the mixing time? Two natural choices are (i) standard, or random scan, Glauber dynamics where the updated variable is chosen uniformly at random, and (ii) the systematic scan dynamics where variables are updated in a fixed, cyclic order. We first show that for systems of dimension $n$, one round of the systematic scan dynamics has spectral gap at most a factor of order $n$ worse than the corresponding spectral gap of a single step of Glauber dynamics, tightening existing bounds in the literature by He, et al. [NeurIPS ’16] and Chlebicka, Łatuszyński, and Miasodejow [Ann. Appl. Probab. ’25]. The corresponding bound on mixing times is sharp even for simple spin systems by an explicit example of Roberts and Rosenthal [Int. J. Statist. Prob. ’15]. We complement this with a converse statement: if all, or even just one scan order rapidly mixes, the Glauber dynamics has a polynomially related mixing time, resolving a question of Chlebicka, Łatuszyński, and Miasodejow.
1 Introduction
@@
An important task in statistics, computer science, and machine learning is sampling from high-dimensional distributions. However, the “curse of dimensionality” can make this problem provably computationally hard in general [Sly10, SS12, GKK24]. Even in settings where sampling may be possible, it is often the case that it is easy to determine the probability of a configuration up to a universal normalizing constant (also called the partition function), but the exact computation of this normalization requires summing over (at least) exponentially many configurations in the dimension. Therefore, a popular option in both theory and practice is to run a suitable ergodic Markov chain on the state space that is easy to implement and whose stationary measure is the desired distribution.
@@
Of these methods, an extremely popular approach for sampling from high-dimensional distributions is the well-known Gibbs sampler introduced by Geman and Geman [GG84] inspired by ideas from statistical physics. The Gibbs sampler is particularly convenient since it is usually quite efficient to implement. Concretely, the general form of these procedures is as follows: for some target distribution $\pi$ on a product space $\mathcal{X}:=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$, consider any starting configuration $X^{0}\in\mathcal{X}$. At each time $t\geq 1$, the sampler chooses a coordinate $i_{t}\in[n]$ according to some predetermined (possibly randomized) rule and rerandomizes the $i_{t}$’th coordinate of $X^{t-1}$ according to the distribution of $\pi$ on $\mathcal{X}_{i_{t}}$ conditioned on the remaining coordinates agreeing with $X^{t-1}$. All other coordinates remain the same, and these conditional distributions are often computationally straightforward to compute. It is not difficult to show that $\pi$ is stationary with respect to these dynamics for any choice of indices $i_{t}$, and thus, the overall complexity of the sampling method is governed by the number of steps for which it needs to run to be close to stationarity.
@@
Two natural ways to choose the coordinate $i_{t}$ to update at time $t$ are (i) standard (random scan) Glauber dynamics, where $i_{t}\in[n]$ is chosen uniformly at random, and (ii) the systematic scan, where $i_{t}=\sigma(t\ (\mathrm{mod}\ n))$ for some fixed choice of permutation $\sigma:[n]\to[n]$ so that the choice of sites to update cycles around the coordinates in some fixed order. For any choice of updates, the crucial quantity of interest is the mixing time to stationarity: how many steps of the Gibbs sampler are needed until the distribution at time $T$ is close to $\pi$ in, say, total variation distance? Ideally, the process mixes in time $\mathsf{poly}(n)$ so that the overall sampling complexity scales only polynomially in the dimension.
The systematic scan is often used in practice due to several implementation advantages over Glauber dynamics [DGJ06b, ZR13]. However, they often lack strong analytic guarantees on the mixing times due to the lack of reversibility. General comparisons between the two chains have been conjectured for several decades [Dia13, LP17], but it is known that the site ordering can in fact substantially affect the mixing time [RR15, HSMR16]: simple examples show that either sampler can dominate the other by $\mathsf{poly}(n)$ factors in mixing time, different scan orders can have a $\mathsf{poly}(n)$ factor effect on the mixing time, and the random scan can lie asymptotically strictly between the fastest and slowest systematic scans. To the best of our knowledge, a precise worst-case comparison of the spectral properties or mixing times of the systematic scan in terms of Glauber has remained open despite significant work comparing them in both special cases and general settings [Ami96, DGJ06a, DR00, BZ23, Sim93, DGJ06b, Fis96, HSMR16, CŁM25]. In the other direction, establishing any explicit polynomial upper bound on the mixing time of Glauber dynamics in terms of the mixing times of the standard systematic scan seems to have remained open as well, appearing as a recent question of Chlebicka, Łatuszyński, and Miasodejow [CŁM25].
In this work, we resolve both of these fundamental questions. In particular, we show the following:
  1. 1.
    We completely resolve the worst-case spectral gap of any systematic scan in terms of the spectral gap of Glauber dynamics: if the Glauber dynamics has spectral gap at least $\gamma$, then the full $n$ steps of the systematic scan has spectral gap at least $\Omega(\gamma/n)$, and this factor is sharp. To the best of our knowledge, this is even the first such comparison result where the ratio between the gaps is independent of the original spectral gap, which can be exponentially small in many known examples where $\gamma=e^{-\Omega(n)}$. Since the scan has $n$ steps, this implies via standard relationships between spectral gaps and mixing times that the number of steps of the scan to mix is at most a factor of $n^{2}\log(1/\pi_{\mathsf{min}})$ larger than that of Glauber. [^1] The logarithmic factor is inherent to spectral methods to analyzing the mixing times; it is an interesting question whether this factor can removed by comparing the modified log-Sobolev constants of these chains. We thus prove that an explicit example of Roberts and Rosenthal [RR15] (namely, the hardcore model on the complete graph) constitutes the provably worst-case instance of mixing times ratio between these chains (up to a constant).
  2. 2.
    Moreover, we establish a polynomial upper bound on the spectral gap of the Glauber dynamics in terms of the spectral gap of systematic scans. By leveraging a recent spectral gap inequality of Chatterjee [Cha25] on the singular values of the Laplacian, this in fact yields a polynomial upper bound in terms of mixing times as well due to the special orthogonality properties of the scan. To the best of our knowledge, this appears to be the first such bound and resolves a question (Open Problem 3.7) of Chlebicka, Łatuszyński, and Miasodejow [CŁM25].
As we describe further below, our analysis naturally extends to any sufficiently pseudorandom sequence of site re-samplings, which may have statistical applications in natural settings where Gibbs sampling is viewed as an exogenous process describing a system’s evolution rather than as a sampling method by a practitioner. An additional bonus of our approach is that it is much more direct, but at the same time sharper, than the significantly more involved functional-analytic approach of Chlebicka, Łatuszyński, and Miasodejow [CŁM25]. Their result, which we describe in more detail in Section 1.3, obtained the previous best, but suboptimal, bound on the spectral gap of the systematic scan that we are aware of. By contrast, our results rely on conceptually natural linear-algebraic and probabilistic arguments that elementary, flexible, and can lead to optimal bounds.
1.1 Background and Our Results
@@
We now describe the background and our results in more detail. By now, a variety of advanced methods have been introduced to bound the mixing time of important Markov chains (see, e.g. [LP17, MT06] for textbook treatments). For high-dimensional distributions on product spaces, mixing time analyses are typically studied for the Glauber dynamics since this Markov chain enjoys several analytical advantages. For one, the Markov chain is reversible with respect to $\pi$ so that the mixing properties can be analyzed by well-understood spectral methods. In particular, the mixing time connects naturally to linear-algebraic machinery via the spectral theorem, making these dynamics significantly more amenable to a rigorous study in most applications. Moreover, the Glauber dynamics treats sites symmetrically and is the natural tensorized analogue of the transition operators that operates independently on each coordinate, as would be the case for product distributions.
@@
By contrast, the transition operator associated with the systematic scan is importantly not typically reversible once $n>2$ except in simple settings, like when $\pi$ is a product measure. Even though the transitions are a product of self-adjoint operators, the lack of commutativity makes the analysis highly nontrivial. In fact, this issue is not just a feature of existing technical machinery: the actual site ordering can indeed significantly affect the mixing time since the application of the transitions varies deterministically depending on the iteration [RR15, HSMR16]. Nonetheless, from the computational side, the fixed and sequential nature of the systematic scan is among several practical considerations that can affect runtime in real-world implementations [ZR13]. No additional randomness is needed in the sequence of updated sites and moreover, one can exploit locality among adjacent sites in the updating process as well as parallelization across distant sites, which can yield practical gains given hardware considerations. For these reasons, experimental evaluations are often done using the systematic scan [DGJ06b].
@@
As a result, it is important to understand the relative performance between these two implementations of the Gibbs sampler. More generally, to what extent does the choice of site updates affect the mixing time of important high-dimensional distributions one wishes to sample from? From an analytical perspective, it may be significantly easier to prove mixing time bounds for Glauber dynamics for the reasons above; but then to what extent can these guarantees extend to the systematic scan? Obtaining (hopefully sharp) black-box guarantees of this type provide concrete performance guarantees for the systematic scan.
@@
In this paper we make significant qualitative and quantitative progress in resolving the worst-case ratios of the mixing times of these two important chains. We first prove the following result on mixing times. To state the result, let $t^{\mathsf{GD}}_{\mathrm{mix}}(\varepsilon)$ denote the time for Glauber dynamics to be within $\varepsilon$ of $\pi$ in total variation distance from any starting configuration, and let $t^{\mathsf{SS}(\sigma)}_{\mathrm{mix}}(\varepsilon)$ denote the time for Glauber dynamics to be within $\varepsilon$ of $\pi$ in total variation distance from any starting configuration, where the scan is done in order $\sigma:[n]\to[n]$ (see Section 2 for more formal definitions).
@@
Since the systematic scan consists of $n$ site updates, Theorem 1.1 asserts that in terms of the number of updates, the mixing time of the systematic scan is no more than a factor $n^{2}$ worse than that of Glauber dynamics up to the logarithmic factor inherent to spectral analyses of mixing. In natural discrete systems, the minimum probability satisfies $\pi_{\min}=2^{-\Theta(n)}$, in which case Theorem 1.1 asserts that the mixing time is at most a factor of $n^{3}$ worse. To the best of our knowledge, this is the first bound showing that the ratio of mixing times is at most polynomial, even in settings where the mixing times are themselves superpolynomially large.
@@
Importantly, Theorem 1.1 is not improvable in the worst-case even in simple (anti-ferromagnetic) spin systems as witnessed by an example of Roberts and Rosenthal [RR15]. For completeness, we provide a simple, self-contained proof of this example following their reasoning in Section 4. Thus, Theorem 1.1 constitutes the sharpest possible upper bound on the mixing time of the systematic scan that can hold unconditionally for all systematic scans for Gibbs samplers in terms of the mixing time of Glauber dynamics, at least up to the unavoidable logarithmic factor.
We in fact show that our spectral gap bound underlying Theorem 1.1 holds for almost every pseudorandom sequence of sites to update. When translated to the Glauber dynamics, or most similar dynamics, our argument yields an efficient certification procedure that a process has mixed, even conditioned on the sites that update, that holds with $1-o(1)$ probability for random uniform sequences. Understanding which deterministic sequences of site updates mix is of importance in applications beyond sampling, like in statistical applications where one does not have perfect knowledge of how sites are chosen to update. For instance, the Glauber dynamics are also considered in economics as “noisy best-response dynamics” for decentralized game dynamics [KMR93, Blu93, You11] as well as in statistical physics as a model for a particle system to converge towards thermal equilibrium. In these systems, one views the dynamics as arising exogenously and instead wishes to learn some macroscopic properties or statistics of typical samples from $\pi$. However, it may not be possible to know the precise (possibly stochastic) process that generates site updates; can the statistician nonetheless conclude that after some predictable period of time, the observed configuration is close to the true stationary measure without knowledge of the precise randomness in the sequence of observed site updates? This amounts to asserting that the configuration has conditional law close to $\pi$ even given the (independently chosen) sequence of updated sites.
Our argument (which we discuss in more detail in Section 1.2) for establishing Theorem 1.1 and these applications turns out to be significantly more flexible than the analysis of Chlebicka, Łatuszyński, and Miasodejow [CŁM25] in that it can be more or less reversed up to polynomial factors. We further show that if any systematic scan has a spectral gap, then Glauber dynamics has a polynomially related spectral gap, which therefore implies a mixing bound by standard arguments. This spectral gap comparison resolves Open Problem 3.7 of Chlebicka, Łatuszyński, and Miasodejow [CŁM25] with the following consequence:
Theorem 1.2.
Let $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$ be a product of finite state spaces and let $\pi$ be any distribution on $\mathcal{X}$. Then for any $0<\varepsilon\leq 1$ and any permutation of $[n]$, $\sigma$,
We provide the precise quantitative dependencies in Section 3; the bound improves depending on the assumption made on whether all systematic scan orders $\sigma$ mix fast, or whether there is a single one that mixes fast. While these bounds are significantly weaker than our tight results for the converse direction, Theorem 1.1 and Theorem 1.2 together constitute an unconditional polynomial equivalence between the mixing times of the Glauber dynamics and any systematic scan. The mixing time comparison follows from the spectral gap comparison by a recent comparison of the mixing time of non-reversible chains in terms of the minimal nontrivial singular value of the Laplacian by Chatterjee [Cha25]. While this value is usually not related to the operator norm for general non-reversible Markov chains, the projection property of the Gibbs sampler will enable us to relate these quantities up to polynomial factors.
1.2 Overview of Techniques
We now turn to giving a high-level overview of our main techniques towards establishing Theorem 1.1 and Theorem 1.2. All precise definitions are given in Section 2.
@@ 1.2 Overview of Techniques
From Glauber Dynamics to Scans. The proof of Theorem 1.1 and the related applications for pseudorandom sequences is based on linear-algebraic arguments that bound the operator norms of appropriate matrices. The connection between spectral gaps and mixing times is fairly standard, so we only briefly remark on this connection here and defer the standard background and proofs to Section 2.
At a high level, it is well-known that the Glauber dynamics having relaxation time to stationarity at most $1/\delta$ is the same as asserting that for any function $f\in L^{2}(\mathcal{X},\pi)$ with $\text{Var}_{\pi}(f)=1$ and $\mathbb{E}_{\pi}[f]=0$,
@@
for appropriate linear operators $\mathsf{P}_{i}$ that are self-adjoint with respect to the natural inner product associated to $L^{2}(\mathcal{X},\pi)$. Here, the norm is the induced operator norm on this space. When acting on distributions on the left, each $\mathsf{P}_{i}$ corresponds to transitioning by rerandomizing in the $i$th site as in the Gibbs sampler. These operators act on functions $f\in L^{2}(\mathcal{X},\pi)$ on the right by taking the conditional expectation of $f$ after rerandomizing the $i$th site given the remaining coordinates, which is well-known to be an orthogonal projection in $L^{2}(\mathcal{X},\pi)$.
Our first main technical result establishes a suitable operator norm bound on the product of transition matrices associated to any systematic scan. Concretely, we prove the following abstract statement:
Theorem 1.3 (Theorem 3.1, restated).
Let $\mathsf{P}_{i_{L}}\ldots\mathsf{P}_{i_{1}}$ be any product of the orthogonal projection operators satisfying the following condition. For each $j\in[n]$, suppose there exists a minimal index $k_{j}\leq L$ satisfying $i_{k_{j}}=j$. We further assume these indices satisfy:
$1=k_{1}\leq k_{2}\leq\ldots\leq k_{n}=L.$ (2)
Then, it holds that
In words, Theorem 1.3 asserts that if the arithmetic mean of orthogonal projection operators satisfies a spectral gap, then the spectral gap of any product of them can be bounded by an explicit quantity that depends only on the first appearances of each distinct projection; this result may be of independent interest. As we discuss in Section 4, this result is completely sharp up to the constant in two respects: first, it is necessary that the operators be orthogonal projections (or else the bound can even be exponentially large!), and second, there are explicit examples of sequences of orthogonal projections [RR12] even in $\mathbb{R}^{2}$ showing the conclusion is tight up to the precise constant.
Returning to the setting of the systematic scan, we have $k_{j}=j$ by definition, in which case the corresponding sum is $\Theta(n^{2})$ (Corollary 3.1). The same can shown with high probability for a sequence of indices drawn uniformly at random from $[n]$ (c.f. Theorem 3.2), and hence a similar contraction holds for any sufficiently pseudorandom sequence. In both cases, the conclusion is that the corresponding spectral gap of the sequence of operators is at least $\Theta(\delta/n)$. By standard arguments relating mixing times and operator norms even for non-reversible Markov chains which we recall in Section 2, we obtain Theorem 1.1. Again, we establish for the first time that a simple example of Roberts and Rosenthal [RR15] yields a tight example for the worst-case mixing comparison.
To establish Theorem 1.3, we use the following reasoning. For each $\ell\geq 1$, we define the operators
$\mathsf{Q}_{\ell}:=\mathsf{P}_{i_{\ell}}\cdots\mathsf{P}_{i_{1}}.$
Our goal is to give an operator norm bound on $\mathsf{Q}_{L}$ and we do so by precisely tracking the spectral gap of each of the intermediate operators $\mathsf{Q}_{\ell}$. The key observation is that for any vector $f\in\mathcal{H}$, we can write the orthogonal decomposition
We exploit this orthogonality with a precise charging argument to bound the norm of the difference $\mathsf{P}_{j}f-f$ in terms of the individual movements $\|\mathsf{Q}_{\ell}f-\mathsf{Q}_{\ell-1}f\|$ up to $\ell=k_{j}$. The intuition is that the spectral gap of Glauber, being an average of such differences, can be charged to the individual movements when we iteratively apply each projection operator. More precisely, if the first application of $\mathsf{P}_{j}$ to $f$ in the iterated projections does not have a similar effect as it does for the average projection, then this must be because the preceding projection operators noticeably moved $f$. But in this case, this increases the spectral gap of the product, since any movement from an orthogonal projection decreases the norm! Making this win-win analysis completely precise follows essentially from the Pythagorean identity that squared norms are exact when written for an orthogonal decomposition. Since the average of operators is assumed to have a spectral gap, we can then ensure that the $\mathsf{Q}_{\ell}$ operators also make progress in decreasing the norm of $f$. In light of Section 4, our reasoning and accounting in this charging argument is completely tight.
From Scans to Glauber. To resolve the polynomial equivalence of the two mixing times, we then turn to establishing Theorem 1.2. A first important subtlety is that for non-reversible chains, the connection between mixing times and spectral gaps is not exact; a spectral gap upper bounds the mixing time, but a lower bound typically only follows from a large norm eigenvalue. In particular, the fact that a Markov chain mixes fast does not in itself imply a good bound on the spectral gap. As a result, it is not immediately clear how to employ similar reasoning to bound the mixing time of Glauber, since this is equivalent to a spectral gap, under just the assumption a scan is fast mixing.
Our first simple contribution, that may be of independent interest, is noting that for the systematic scan, there is such a lower bound. In particular, we exploit a recent connection of Chatterjee [Cha25] that establishes lower bounds on mixing times in terms of spectral gaps of the Laplacian of the scan. For a general Markov chain $\mathsf{P}$, the Laplacian is defined as $\mathsf{I}-\mathsf{P}$, and there is not usually a direct comparison between the singular values of the Laplacian and the original chain when nonreversible. However, we show in Lemma 2.1 that there is such a comparison for the scan precisely due to the fact each individual operator in the systematic scan is an orthogonal projection.
With this in order, we can then aim to establish a converse on bounding the mixing time of Glauber via one or all scans in terms of spectral gaps. The key technical claim we show is the following:
Theorem 1.4 (Theorem 3.4, restated).
$1=k_{1}\leq k_{2}\leq\ldots\leq k_{n}=L.$ (4)
Then, it holds that
In words, we show that if a fixed sequence of the orthogonal projections has a spectral gap, then the same will be true of any supersequence where the spacing between the first appearances of the operators in order is not too large. This statement and conclusions are qualitatively similar to Theorem 1.3, but loses in the square due to less favorable orthogonality arguments that are lossy. To prove Theorem 1.4, we employ a similar careful win-win analysis; either the first application of a projection $\mathsf{P}_{j}$ affects a given vector in similar ways, or the loss can be charged to the additional projections in the supersequence. Balancing these contributions properly leads to the final bound.
To apply this result to bounding the spectral gap of Glauber, we use the following key observation: since the Glauber dynamics is reversible, the operator is self-adjoint, and hence
$\left\|\frac{1}{n}\sum_{i=1}^{n}\mathsf{P}_{i}\right\|^{L}=\left\|\left(\frac{1}{n}\sum_{i=1}^{n}\mathsf{P}_{i}\right)^{L}\right\|.$
Therefore, it suffices to bound the spectral gap for a suitable power. The key idea is that taking $L\approx n\log(n)$ or $L\approx n^{2}\log(n)$, expanding the power leads to sums of products of the projections that with very high probability contain at least one scan or some fixed scan ordering as an embedded subsequence. In particular, we may apply Theorem 1.4 to assert that most of these terms have an explicit spectral gap, while we may trivially bound the rest of the terms by $1$. Since the probability of a random sequence of $L$ projections containing a scan decays much much faster than $L$, the loss in the denominator of Theorem 1.4 is much less than the error probability of not containing a good scan. Setting parameters precisely leads to the polynomial equivalence of Theorem 1.2, thus resolving the open problem of Chlebicka, Łatuszyński, and Miasodejow [CŁM25].
@@
Organization. In Section 1.3, we discuss related bounds and comparisons in the literature between Glauber dynamics and the systematic scan; to our knowledge, our spectral gap comparison subsumes all such existing bounds and cannot be improved in light of known lower bounds. In Section 2, we include for completeness standard results on the relation between the spectral gap and mixing times, including for non-reversible chains, for our main application. In Section 3, we prove the requisite linear-algebraic facts that yield our mixing time comparisons. Finally, we show in Section 4 that Theorem 1.1 and Theorem 1.3 cannot be significantly improved via simple, existing examples.